OpenStudy (anonymous):

What are the solutions of 4x2 + x = –3?

OpenStudy (anonymous):

\[4x^2+x=-3\] i guess You have to solve that equation, you'll obtain: \[4x^2+x+3=0\] using the "formula" to obtain the roots you'll obtain \[\frac{ 1 }{ 8 } \pm i \sqrt{47}\]

OpenStudy (anonymous):

Perfect, thank you!